The motion of a certain spring-mass system is governed by the differential equation d^2u/dt^2+1/8du/dt+u=0 where u is measured in metres and t in seconds. If u(0)=2 and (du/dt)t=0=0 determine the subsequent motion. Also determine the time at which the mass first passes through the equilibrium position.
1
Expert's answer
2020-12-04T10:54:56-0500

Let us solve the equation:


"\frac{d^2u}{dt^2}+\frac{1}{8}\frac{du}{dt}+u=0"


The characteeristic equation "k^2+\frac{1}{8}k+1=0" is equivalent to "(k+\frac{1}{16})^2=\frac{1}{256}-1=-\frac{255}{256}"


So, "k+\frac{1}{16}=\pm i\frac{\sqrt{255}}{16}" and thus "k=\frac{-1\pm i\sqrt{255}}{16}".


Therefore, the solution of the differential equation is


"u=e^{-\frac{t}{16}}(C_1\cos(\frac{\sqrt{255}}{16} t)+C_2\sin(\frac{\sqrt{255}}{16} t))"


Then "\frac{du}{dt}=-\frac{1}{16}e^{-\frac{t}{16}}(C_1\cos(\frac{\sqrt{255}}{16} t)+C_2\sin(\frac{\sqrt{255}}{16} t))+e^{-\frac{t}{16}}(-C_1\frac{\sqrt{255}}{16}\sin(\frac{\sqrt{255}}{16} t)+C_2\frac{\sqrt{255}}{16}\cos(\frac{\sqrt{255}}{16} t))"


If "u(0)=2" and "\frac{du}{dt}(0)=0" then we have "2=u(0)=C_1" and "0=\frac{du}{dt}(0)=-\frac{1}{16}C_1+C_2\frac{\sqrt{255}}{16}" .


Therefore, "C_1=2" and "C_2=\frac{2}{\sqrt{255}}" and the subsequent motion is the following:


"u=e^{-\frac{t}{16}}(2\cos(\frac{\sqrt{255}}{16} t)+\frac{2}{\sqrt{255}}\sin(\frac{\sqrt{255}}{16} t))=
2e^{-\frac{t}{16}}(\cos(\frac{\sqrt{255}}{16} t)+\frac{1}{\sqrt{255}}\sin(\frac{\sqrt{255}}{16} t))."


Let us determine the time at which the mass first passes through the equilibrium position. For this we should solve the equation:


"u(t)=0"


"2e^{-\frac{t}{16}}(\cos(\frac{\sqrt{255}}{16} t)+\frac{1}{\sqrt{255}}\sin(\frac{\sqrt{255}}{16} t))=0"


"\cos(\frac{\sqrt{255}}{16} t)+\frac{1}{\sqrt{255}}\sin(\frac{\sqrt{255}}{16} t)=0"


"\tan(\frac{\sqrt{255}}{16} t)=-\sqrt{255}"


"\frac{\sqrt{255}}{16} t\approx -1.51+3.14 n, \ n\in\mathbb Z."


Since the mass first passes through the equilibrium position and "t\ge 0", we conclude that "n=1". Therefore, "\frac{\sqrt{255}}{16} t\approx -1.51+3.14=1.63" and "t\approx 1.63\frac{16}{\sqrt{255}}\approx 1.633" (seconds).